OpenStudy (anonymous):

Please help will fan and give medal The function f(t) = 30 sin (pi over 3t) −15 models the temperature of a periodic chemical reaction where t represents time in hours. What are the maximum and minimum temperatures of the reaction, and how long does the entire cycle take? maximum: 30°; minimum: 15°; period: pi over 3 hours maximum: 30°; minimum: −15°; period: pi over 3 hours maximum: 45°; minimum: 15°; period: 6 hours maximum: 15°; minimum: −45°; period: 6 hours

OpenStudy (anonymous):

@ganeshie8 @amistre64 @tanmaykhandait

OpenStudy (anonymous):

@jacobwat

OpenStudy (anonymous):

Sin range fro [-1,1] Therefore 30sin theta ranges ranges from [-30,30] Subtracting 15 The function ranges from from [-45,15]

OpenStudy (anonymous):

so @tanmaykhandait would you say d? because based on your info thats what im thinking

OpenStudy (anonymous):

Ya

OpenStudy (anonymous):

awesome can you help me witha few more questions @tanmaykhandait

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

awesome thank you :-) im typing it now @tanmaykhandait

OpenStudy (anonymous):

@tanmaykhandait the function f(t)=5 cos (pi/4t)+11 represents the tide in dark sea. it has a maximum of 16 feet when t time t is 0 and a minimum of 6 feet. the sea repeats this cycle every 8 hours after 6 hours what is the tide? a.)11 b.)6 c.)8.5 d.)10.5

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Sorry...I don't know...will let u know in some time

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @helpmemath123 @tanmaykhandait the function f(t)=5 cos (pi/4t)+11 represents the tide in dark sea. it has a maximum of 16 feet when t time t is 0 and a minimum of 6 feet. the sea repeats this cycle every 8 hours after 6 hours what is the tide? a.)11 b.)6 c.)8.5 d.)10.5 \(\color{blue}{\text{End of Quote}}\) This is just a matter of plugging in \(t=6\): \[f(6)=5\cos\frac{6\pi}{4}+11=0+11\]

OpenStudy (anonymous):

thank you this is the only one i got wrong on my exam :)